For the right rectangular prism (or box) shown, the dimensions are = 12 in, w = 4 in,and h = 3 in. How long is a diagonal of the box?
Answer by stanbon(75887)

For the right rectangular prism (or box) shown, the dimensions are = 12 in, w = 4 in,and h = 3 in. How long is a diagonal of the box?
-------
diagonal of the base = sqrt[12^2+4^2] = sqrt(160)
-------
Diagonal of the box = sqrt[160 + 3^2] = sqrt(169) = 13 inches
============================================================
Cheers,
Stan H.
============